OpenStudy (anonymous):

please help! a equation of the line that contains the origin and the point (1,2) is A. y=2x B. 4 C. 5 D. 6 E.7

OpenStudy (phi):

first, find the slope between (0,0) and (1,2) use change in y divided by change in x

OpenStudy (anonymous):

2/-1?

OpenStudy (anonymous):

So its A?

OpenStudy (phi):

(2-0)/(1-0) = 2/1 = 2 so the slope is 2 y = 2x + b use one of the points (0,0) is the easiest replace y with 0 and x with 0 we get 0= 2*0 + b 0= b so b is zero, and y =2x+ b becomes y=2x

OpenStudy (anonymous):

ok thanks @phi :)